jennylove:

In the circuit shown, a current of 0.33 A will pass through which of the following resistors? R3 R1 R2 R4

jennylove:

1 attachment
jennylove:

R1?

jennylove:

@justjm if you could help as well?

TheSmartOne:

I've been looking at this far too long and I'm stumped Are you sure that the question/image are correct? I'm not able to get 0.33 at all mhmm

TheSmartOne:

Or if you could show how you got to that and I'm just completely overthinking it

jennylove:

actually nvm that r1 would be for 6.6 V because (0.78)(8.6) = 6.6

jennylove:

im not sure 0.33 :/

TheSmartOne:

Are you sure you're not misreading it and we're supposed to find 0.83 and not 0.33

jennylove:

R1 and R4 and the parallel combo R2 R3 all have the same current through them (I did not calculate it......it will likely be r2 or r3 since you cannot choose multiple )

jennylove:

right?

jennylove:

idk

jennylove:

and no it is 0.33, its not an error :/

TheSmartOne:

Yes you're right So for parallel, the voltage is the same and the resistance would be 1/(1/13 + 1/18) = 7.55 ohms And voltage is 15 V (stays same in parallel, changes in series) So 7.55 + 8.5 + 3.2 = 19.25 ohms V = IR So current for the resistors in series will all be the same which would be 15/19.25 = 0.779 A So it can't be resistor 1 or 4

TheSmartOne:

I think we're not considering the voltage drop after resistor 1

jennylove:

hmm , okay yes that does make some more sense.. im honestly not sure

TheSmartOne:

@imqwerty whenever you get a chance and you're free, if you could take a look at this. We could use your intelligence:))

jennylove:

would it be okay if i closed this question, just so i can upload more? until @imqwerty esponds

jennylove:

responds

TheSmartOne:

Yea, for sure

TheSmartOne:

Circuits and all are always a little tricky >,<

jennylove:

very true! ive been having a lot of difficulty

imqwerty:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TheSmartOne Yes you're right So for parallel, the voltage is the same and the resistance would be 1/(1/13 + 1/18) = 7.55 ohms And voltage is 15 V (stays same in parallel, changes in series) So 7.55 + 8.5 + 3.2 = 19.25 ohms V = IR So current for the resistors in series will all be the same which would be 15/19.25 = 0.779 A So it can't be resistor 1 or 4 \(\color{#0cbb34}{\text{End of Quote}}\) this is correct, it can't the resistor 1 or 4, so we need to check the current in resistors 3 and 2

imqwerty:

a total current of 0.779Amps flows in the circuit

imqwerty:

When the current arrives at the junction connecting wire of resistors 3 and 2, it splits up. So a part of the 0.779 A goes into resistor 2 and the remaining goes into resistor 3

imqwerty:

how to calculate the amount of current that goes into each resistor? ez- the more the resistance the lesser the current. current in r2 = \(\large {I \times \frac{r_2}{r_2+r_3}}\) current in r3 = \(\large {I \times \frac{r_3}{r_2+r_3}}\) where I is the total current arriving at the junction = 0.799 A

imqwerty:

how did I get that formula to find the currents in each branch? - the potential drop across r2 and r3 will be the same because they have the same endpoints assume that the potential drop is \(V\) so \(V = r_2 \times i_2\) and \(V = r_3 \times i_3\) also \(i_2 + i_3 = I \) where \(i_2\) = current in resistor r2 and \(i_3 \) = current in resistor r3 solving the above equations you can get the expression for \(i_2 \) and \(i_3\) that i wrote in my previous message

TheSmartOne:

you're a genius qwerts <3

imqwerty:

master <3